# Depth-guided NeRF Training via Earth Mover’s Distance

### v1

@@ 4.4 Role of Depth Supervision and Uncertainty
We wish to better understand how NeRFs leverage depth priors during training. Thus, to examine the role of depth prior and depth uncertainty in downstream performance, we conduct extensive ablations studies under various settings.
In Table 3 (a) we evaluate the importance of design choices of our model compared to a basline NeRF model. We ablate Regularization $Reg.$, use of depth supervision $z$ (by default with L2 loss), use of $EMD$ depth supervision, and uncertainty weighting $u$. As we can observe, depth supervision plays a vital role in improving results, especially depth metrics. But depth alone ($Reg.$ + $z$) does not yield satisfactory depth results. Replacing a naive $L_{2}$-loss with the EMD ($Reg.$ + $z$ + EMD + $u$) considerable reduces the depth error, highlighting the importance of depth guidance.
Table 3: Ablation studies for the components of our method, varying depth priors, and varying objectives for leveraging these priors. All reported results are on ScanNet scenes.
[…]
Although the impact of $u$ is small when averaged across entire images, $u$ makes a difference where it matters: in uncertain areas. In Figure 7 (a), we first evaluate how well the uncertainty measure can distinguish reliable from unreliable pixels. We compute the depth error of DiffDP for those pixels that have an uncertainty of $\geq t$ and compare it to more certain pixels with $u<t$ for thresholds $t=\{0,0.1,...0.8\}$. For a threshold of, for instance, $t=0.5$, the depth error of uncertain pixels is $63\%$ higher than the error of certain pixels. Our uncertainty measure is inherent to the depth generation process and displays an impressive capability to highlight pixels that should not be trusted.
[…]
@@
To evaluate its impact on our model, we ablate the use of $u$ in Figure 7 (b) in a similar manner as the ablation in Table 3 (a); but instead of averaging over an entire image, we look at uncertain and certain regions of an image. We observe, that most pixels have a low uncertainty (64% of all pixels have an uncertainty of $<0.1$), and for those pixels incorporating uncertainty only marginally (1.9%) helps. But for pixels with high uncertainty (>0.8), incorporating $u$ improves the RMSE by 5.1%.
[…]
[width=]figures/psnr_depth_cimle.eps
Figure 8: Multimodality implies degraded performance. The depth error increases rapidly when arranged in ascending order.
Figure 9: We repeat the experiment in Table 1 with four random seeds and show the depth RMSE and PSNR of each of the three scenes in ScanNet. While SCADE can sometimes produce accurate depths, the variability between runs is large. Our method produces consistently accurate depths.
@@
It is important to note that our model’s ability to predict accurate depth does not solely stem from a good prior. We compare our prior to SCADE’s LeReS-based prior [31] in Figure 9, where we adjusted the scale between prior and ground truth as ratio between their mean depths before calculating the RMSE. We make two interesting observations. First, sorting SCADE’s 20 priors by accuracy, we find that the error increases rapidly. The additional priors introduce noise rather than adding beneficial information. This might explain why SCADE’s training is at times unstable, and why learning accurate depth is so important for NeRFs. In fact, we can observe in Figure 9 that SCADE depends highly on the chosen random seed. The depth RMSE varies dramatically between runs. Comparing this to the PSNR, we observe that even when the depth prediction is widely off, the NeRF can still produce decent RGB quality with a PSNR of at least 20 for scene 781. However, the model learned a wrong underlying scene, as the maximum depth error of 1.6 illustrates. So while a NeRF can still learn to produce realistic looking RGB images, it has no knowledge of the underlying geometry (as seen in Figure 4). This not only prevents applications such as augmented reality in which a user would interact with the 3D world or when meshes need to be extracted for the scene, but also renders reported photometric results unreliable.
@@
In Table 3 (b), we evaluate the initial strength of our DiffDP based prior and the prior from [29]. We show that DiffDP is a good choice for depth prior even in comparison to priors generated from large-scale pretraining like in DepthAnything [29]. Given this context, Table 3 (c) exemplifies our $EMD$-based objective’s role in leveraging information from the priors. Irrespective of prior, our loss encourages a reduction in the initial prior’s depth error (with our DiffDP prior, RMSE drops 31.4%, and with DepthAnything prior, RMSE drops 18.8%). Through $EMD$ depth guidance, NeRFs can leverage these priors without enforcing them, leading to a smaller overall depth error. To further examine the impact of our objective, we construct a strong baseline under different depth priors that mimics the loss presented in [25] but with a single depth hypothesis. Specifically, we use L2 loss on depth hypotheses during NeRF training. We make two observations in Table 3 (d): First, depth metrics remain constant or even worsen in comparison to the initial depth priors from (b) under the L2 objective. Secondly, our $EMD$-loss strongly outperforms the L2 objective for both priors. These experiments highlight the importance of supervising NeRF with an EMD-based objective that can selectively incorporate information from the depth prior irrespective of the prior’s strength.
@@ Appendix 0.A Additional Qualitative Results
[…]
[width=]app_figures/qualitative_results_depth_errs_781.eps
@@
[…]
[width=trim=0 10.5cm 0 0,clip]app_figures/uncertainty_overview_both.eps
@@
To supplement Section 4.3 in the manuscript, we report additional qualitative results. Figure S1 depicts depth predictions and corresponding error maps for all test images in one of ScanNet’s scenes [5]. Similar to the results in Figure 5, we observe more artifacts in SCADE’s [25] predictions. DäRF’s [24] depth looks promising but leads to a high depth error, manifesting in a brighter red color of the error maps compared to our method. These findings are consistent with the quantitative results in Table 2.
@@ Appendix 0.B Additional Ablation Studies
[…]
0.75! PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$ Ours w/ DiffDP (original) 21.69 0.737 0.373 0.070 0.221 SCADE w/ DiffDP 21.51 0.732 0.376 0.091 0.319 SCADE w/ LeRes cIMLE (original) 21.54 0.732 0.292 0.086 0.252 DäRF w/ DiffDP 21.19 0.762 0.324 0.086 0.279 DäRF w/ MiDaS (original, retrained) 21.55 0.766 0.319 0.085 0.259
@@
[…]
! RGB-based metrics Depth-based metrics PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ SqRel $\downarrow$ RMSE $\downarrow$ RMSE log $\downarrow$ Ours w/ LeRes cIMLE 20.23 0.654 0.457 0.097 0.159 0.915 0.194 SCADE* w/ LeRes cIMLE (original) 19.83 0.664 0.347 0.098 0.171 0.937 0.213
@@
Conversely, we investigated whether our framework can leverage SCADE’s prior in indoor environments. We are especially interested in out-of-distribution scenes to evaluate situations in which the depth prior is uncertain. In Table S2 we compare the use of the LeRes cIMLE prior [25] in both SCADE and our method on the indoor Tanks and Temples scenes used by SCADE. We can observe that our method produces smaller depth errors using the same priors as SCADE. As the Tanks and Temples scenes differ from standard scenes used to pretrained the depth prior (a church, a courtroom, and an auditorium), we hypothesize that our method leveraging Earth Mover’s Distance is better equipped to capture the uncertainty in the depth priors. Note that during training we use all 20 depth hypothesis to compute the $EMD$ loss.
@@ Appendix 0.C Outdoor Experiments
Although our DiffDP prior was trained on indoor scenes, while DäRF’s MiDaS prior was trained on indoor and outdoor data, our method outperforms DäRF on both depth metrics in Table S3. To fully elucidate the influence of priors versus the model design, we additionally trained both methods using the same DepthAnything depth priors obtained from the official outdoor metric weights [29]. We can observe that even when both models leverage the same prior, our model outperforms DäRF on the depth metrics. Interestingly, our prior, although trained indoor, leads to the best geometric results. One reason might be that DiffDP with our uncertainty measure can compensate for uncertainty in outdoor environments, such as sky or far away surfaces. We show qualitative results for both scenes in Figures S3 and S4. Similar to the results on ScanNet in Figure 6 of the main paper, DäRF’s depth maps look crisp. This does not speak to the accuracy of the rendered depth but to the strict enforcement of the MiDaS prior. Rendered depths may look well while leading to high errors (see also Figure S1 for further examples of this behavior).
@@
[…]
0.75! PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$ Ours w/ DiffDP (original) 16.39 0.462 0.692 0.216 1.539 Ours with DepthAnything 16.05 0.457 0.701 0.248 2.193 DäRF w/ MiDaS (original) 16.14 0.497 0.668 0.580 2.362 DäRF w/ DepthAnything 16.14 0.499 0.669 0.579 2.361
[…]
[width=]app_figures/qualitative_results_depth_additionaltruck.eps
@@
[…]
[width=]app_figures/qualitative_results_depth_Ignatius.eps
@@ 0.D.1 Our method
[…]
@@ 0.D.2 DäRF details
We use the rescaled depth maps (better than author-reported) for comparisons in Figures S1 and S2. For the results in Table 2, we assume a scale of 1, since the authors do not report details on how they obtained scales used in their code base.

### v2

@@ 4.4 Role of Depth Supervision and Uncertainty
We wish to better understand how NeRFs leverage depth priors during training and conduct extensive ablations studies under various settings.
In Table 3 (a) we evaluate the importance of design choices of our model compared to a basline NeRF model. We ablate Regularization $Reg.$, use of depth supervision $z$ (by default with $L_{2}$ loss), use of $EMD$ depth supervision instead of $L_{2}$ loss, and uncertainty weighting $u$. We observe that depth supervision plays a vital role in improving results, especially depth metrics. But depth alone ($Reg.$ + $z$) does not yield satisfactory depth results. Replacing a naive $L_{2}$-loss with the EMD ($Reg.$ + $z$ + EMD + $u$) considerable reduces the depth error, highlighting the importance of depth guidance.
Table 3: Ablation studies for the components of our method, varying depth priors and the objectives leveraging them. All reported results are on ScanNet scenes.
Reg. $z$ $EMD$ $u$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$
21.28 0.735 0.370 0.119 0.354
21.24 0.734 0.377 0.100 0.295
21.70 0.736 0.369 0.071 0.222
21.69 0.737 0.377 0.070 0.221
(a) Impact of Regularization $Reg.$, the DiffDP prior $z$, the Earth Mover’s distance loss $EMD$, and uncertainty $u$ on the baseline NeRF model.
AbsRel $\downarrow$ SqRel $\downarrow$ RMSE $\downarrow$ RMSE log $\downarrow$
DiffDP 0.125 0.056 0.324 0.151
DepthAnything 0.144 0.061 0.357 0.180
(b) Comparison of our depth prior DiffDP [11] and DepthAnything [30] evaluated on the training images used as input to our model.
With our loss ($EMD$) PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE $\downarrow$
Ours w/ DiffDP‡ 21.70 0.736 0.369 0.222
Ours w/ DepthAnything‡ 21.62 0.732 0.378 0.290
(c) Impact of the depth priors in (b) on our model. For both priors, our $EMD$ objective allows the NeRF to learn smaller depth errors than the priors used for supervision yield.
With $L_{2}$-H loss PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE $\downarrow$
Ours w/ DiffDP‡ 21.34 0.726 0.382 0.334
Ours w/ DepthAnything‡ 21.04 0.717 0.394 0.349
(d) Comparison of different depth priors under a $L_{2}$-objective applied to depth hypotheses ($L_{2}$-H loss $\equiv$ space-carving loss [26] with only one depth hypothesis). ‡Without uncertainty $u$, as it cannot be used with DepthAnything prior.
Although the impact of $u$ is small when averaged across entire images, $u$ makes a difference where it matters: in uncertain areas. In Figure 7 (a), we evaluate how well the uncertainty measure can distinguish reliable from unreliable pixels. We compute the depth error of DiffDP for pixels that have an uncertainty of $\geq t$ and compare it to more certain pixels with $u<t$ for thresholds $t=\{0,0.1,...0.8\}$. For $t=0.5$, for instance, the depth error of uncertain pixels is $63\%$ higher than the error of certain pixels. Our uncertainty measure thus reliably highlights pixels that should not be trusted. To evaluate its impact on our model, we ablate the use of $u$ in Figure 7 (b) in a similar manner as the ablation in Table 3 (a); but instead of averaging over an entire image, we look at uncertain and certain regions of an image. We observe, that most pixels have a low uncertainty (64% of all pixels have an uncertainty of $<0.1$), and for those pixels incorporating uncertainty only marginally (1.9%) helps. But for pixels with high uncertainty (>0.8), incorporating $u$ improves the RMSE by 5.1%.
@@
It is important to note that our model’s ability to predict accurate depth does not solely stem from a good prior. We compare our prior to SCADE’s LeReS-based prior [32] in Figure 9, where we adjusted the scale between prior and ground truth as ratio between their mean depths before calculating the RMSE. We make two interesting observations. First, sorting SCADE’s 20 priors by accuracy, we find that the error increases rapidly. The additional priors introduce noise rather than adding beneficial information. This might explain why SCADE’s training is at times unstable, and why learning accurate depth is so important for NeRFs. We can observe in Figure 9 that SCADE depends highly on the chosen random seed. The depth RMSE varies dramatically between runs. Comparing this to the PSNR, we observe that even when the depth prediction is widely off (depth error of 1.6 in scene 781), the NeRF can still produce decent RGB quality with a PSNR of at least 20. So while a NeRF can learn to produce realistic looking RGB images, it has no knowledge of the geometry (as seen in Figure 4). This not only prevents applications such as augmented reality in which a user would interact with the 3D world or when meshes need to be extracted for the scene, but also renders reported photometric results unreliable.
Figure 8: Depth errors of cIMLE priors increase rapidly when arranged in ascending order.
Figure 9: We repeat the experiment in Table 1 with four random seeds and show the depth RMSE, PSNR, and LPIPS for each scene in ScanNet. While SCADE can sometimes produce accurate depths (low RMSE), the variability between runs is large. Our method produces consistently accurate depths.
@@
In Table 3 (b), we evaluate the initial accuracy of our DiffDP based prior and the prior from [30]. We show that DiffDP is a good choice for depth prior even in comparison to priors generated from large-scale pretraining like in DepthAnything [30]. Given this context, Table 3 (c) exemplifies our $EMD$-based objective’s role in leveraging information from the priors. Irrespective of prior, our loss encourages a reduction in the initial prior’s depth error (with our DiffDP prior, RMSE drops 31.4%, and with DepthAnything prior, RMSE drops 18.8%). Through $EMD$ depth guidance, NeRFs can leverage these priors without enforcing them, leading to a smaller overall depth error. To further examine the impact of our objective, we construct a strong baseline under different depth priors that mimics the loss presented in [26] but with a single depth hypothesis. Specifically, we use $L_{2}$ loss on depth hypotheses during NeRF training. We make two observations in Table 3 (d): First, depth metrics remain constant or even worsen in comparison to the initial depth priors from (b) under the $L_{2}$ objective. Secondly, our $EMD$-loss strongly outperforms the $L_{2}$ objective for both priors. These experiments highlight the importance of supervising NeRF with an EMD-based objective that can selectively incorporate information from the depth prior irrespective of the prior’s strength.
@@ 5 Conclusions
Acknowledgements
Thanks to the anonymous reviewers for their constructive feedback. This work was supported by the Isackson Family Foundation, the Stanford Head and Neck Surgery Research Fund, and the Stanford Graduate Fellowship.
@@ Appendix 0.A Additional Qualitative Results
Figure S1: Comparison of the learned geometry in 3D. Each plot encompasses the concatenated 3D clouds of two test images with small overlap of each scene in ScanNet. Depth and RGB values are rendered. Error maps in m. Invalid ground truth values are set to zero (yellow).
To supplement Section 4.3 in the manuscript, we report additional qualitative results. We project the learned geometry and the rendered RGB values into 3D in Figure S1. In each example, we show two test images from a scene concatenated into a mutual 3D cloud with some but minimal overlap between the images. We also project the depth errors onto the 3D structure, where darker colors indicate higher errors.
@@
Figure S2 depicts depth predictions and corresponding error maps for all test images in one of ScanNet’s scenes [6]. Similar to the results in Figure 5, we observe more artifacts in SCADE’s [26] predictions. DäRF’s [25] depth looks promising but leads to a high depth error, manifesting in a brighter red color of the error maps compared to our method. These findings are consistent with the quantitative results in Table 1.
@@ Appendix 0.B Additional Ablation Studies
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$
Ours w/ DiffDP (original) 21.69 0.737 0.373 0.070 0.221
SCADE w/ DiffDP 21.51 0.732 0.376 0.091 0.319
SCADE w/ LeRes cIMLE (original) 21.54 0.732 0.292 0.086 0.252
DäRF w/ DiffDP 21.19 0.762 0.324 0.086 0.279
DäRF w/ MiDaS (original, retrained) 21.55 0.766 0.319 0.085 0.259
@@
RGB-based metrics Depth-based metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ SqRel $\downarrow$ RMSE $\downarrow$ RMSE log $\downarrow$
Ours w/ LeRes cIMLE 20.23 0.654 0.457 0.097 0.159 0.915 0.194
SCADE* w/ LeRes cIMLE (original) 19.83 0.664 0.347 0.098 0.171 0.937 0.213
@@
Conversely, we investigated whether our framework can leverage SCADE’s prior in indoor environments. We are especially interested in out-of-distribution scenes to evaluate situations in which the depth prior is uncertain. In Table S2 we compare the use of the LeRes cIMLE prior [26] in both SCADE and our method on the indoor Tanks and Temples scenes used by SCADE. Some qualitative results are shown in Figure S4. For all Tanks and Temples experiments, we initialize the learnable scale parameter as mean over all ratios between sparse SfM estimates and depth prior as proposed by SCADE. We can observe that our method produces smaller depth errors using the same priors as SCADE. As the Tanks and Temples scenes differ from standard scenes used to pretrained the depth prior (a church, a courtroom, and an auditorium), we hypothesize that our method leveraging Earth Mover’s Distance could be better equipped to capture the uncertainty in the depth priors. However, these results should be interpreted with care, as we only have information from very sparse SfM ground truth typically located around edges. Note that during training we use all 20 depth hypothesis to compute the $EMD$ loss.
Figure S4: Qualitative results on indoor TnT scene Church. Note that some geometrical details are better preserved by our method (green box), although the same depth prior is used.
Lastly, for completeness, in addition to comparing the depth priors DiffDP and DepthAnything under $EMD$ and $L_{2}$-H losses in Tables 3(c) and 3(d), we compare the priors under the standard $L_{2}$ loss in Table S3. Unsurprisingly, DiffDP, the more accurate prior (see Table 3(a)) leads to a smaller depth error.
Table S3: Loss ablation study for depth priors DiffDP and DepthAnything. In addition to $EMD$ and space carving loss in Table 3(c,d), we add results using $L_{2}$ loss.
$L_{2}$ loss PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE $\downarrow$
Ours w/ DiffDP‡ 21.24 0.734 0.377 0.295
Ours w/ DepthAnything‡ 21.14 0.737 0.364 0.353
‡Without uncertainty $u$, as it cannot be used with DepthAnything prior.
@@ Appendix 0.C Outdoor Experiments
Although our DiffDP prior was trained on indoor scenes, while DäRF’s MiDaS prior was trained on indoor and outdoor data, our method outperforms DäRF on both depth metrics in Table S4. To fully elucidate the influence of priors versus the model design, we additionally trained both methods using the same DepthAnything depth priors obtained from the official outdoor metric weights [30]. We can observe that even when both models leverage the same prior, our model outperforms DäRF on the depth metrics. Interestingly, our prior, although trained indoor, leads to the best geometric results. One reason might be that DiffDP with our uncertainty measure can compensate for uncertainty in outdoor environments, such as sky or far away surfaces. We show qualitative results for both scenes in Figures S5 and S6. Similar to the results on ScanNet in Figure 6 of the main paper, DäRF’s depth maps look crisp. This does not speak to the accuracy of the rendered depth but to the strict enforcement of the MiDaS prior. Rendered depths may look well while leading to high errors (see also Figure S2 for further examples of this behavior).
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$
Ours w/ DiffDP (original) 16.39 0.462 0.692 0.216 1.539
Ours with DepthAnything 16.05 0.457 0.701 0.248 2.193
DäRF w/ MiDaS (original) 16.14 0.497 0.668 0.580 2.362
DäRF w/ DepthAnything 16.14 0.499 0.669 0.579 2.361
@@ 0.D.1 Our method
$\tau$ $=(\text{max}\textunderscore\text{depth}-\text{min}\textunderscore\text{depth})\cdot 0.0001,$ (9)
@@ 0.D.2 DäRF details
We use the rescaled depth maps (better than author-reported) for comparisons in Figures S2 and S3. For the results in Table 2, we assume a scale of 1, since the authors do not report details on how they obtained scales used in their code base.